# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v1

@@ Abstract
Bacteria evolve in volatile environments and complex spatial structures. Migration, fluctuations and environmental variability therefore have a significant impact on the evolution of microbial populations. Here, we consider a class of spatially explicit metapopulation models arranged as regular (circulation) graphs where wild-type and mutant cells compete in a time-fluctuating environment where demes (subpopulations) are connected by slow cell migration. The carrying capacity is the same at each deme and endlessly switches between two values associated to harsh and mild environmental conditions. When the rate of switching is neither too slow nor too fast, the dynamics is characterised by bottlenecks and the population is prone to fluctuations or extinction. We analyse how slow migration, spatial structure, and fluctuations affect the phenomena of fixation and extinction on clique, cycle, and square lattice metapopulations. When the carrying capacity remains large, bottlenecks are weak and deme extinction can be ignored. The dynamics is thus captured by a coarse-grained description within which the probability and mean time of fixation are obtained analytically. This allows us to show that, in contrast to what happens in static environments, the mutant fixation probability depends on the rate of migration. We also show that the fixation probability and mean fixation time can exhibit a non-monotonic dependence on the switching rate. When the carrying capacity is small under harsh conditions, bottlenecks are strong, and the metapopulation evolution is shaped by the coupling of deme extinction and strain competition. This yields rich dynamical scenarios, among which we identify the best conditions to eradicate mutants without dooming the metapopulation to extinction. We offer an interpretation of these findings in the context of an idealised treatment and discuss possible generalisations of our models.
@@ I Introduction
In the next section, we introduce the explicit spatial metapopulation model that we study and outline our methodology. In Sec. III, we present our results in the case of static environments. This paves the way to the detailed analysis in time-fluctuating environments of Sec. IV, with the weak and strong bottleneck regimes respectively studied in Secs. IV.0.1 and IV.0.2. Sec. V is dedicated to a discussion of our findings, assumptions and possible generalisations. We present our conclusions in Sec. VI. Additional technical details are given in a series of appendices.
@@ II Model & Methods
Figure 1: Metapopulation dynamics in a static environment. (a) Examples of metapopulation graphs: a clique, cycle, and grid (from left to right). Neighbouring demes are connected by migration (double arrows). Initially, there is one mutant deme (red/light) and $\Omega-1$ wild-type demes (blue / dark), and all demes have the same constant carrying capacity $K$. (b) Dynamics in a single deme. Left: Wild-type $W$ cells (blue / dark) compete with mutants of type $M$ (red / light). When $K$ is small, the deme is prone to extinction. When $K$ is large, both types coexist prior to $W$ or $M$ fixation. Top, right: Several realisations of the rescaled deme size $n/K$ vs. time $t$ for $K=5$ (orange/light) and $K=100$ (green/dark). Bottom, right: Fraction of $M$ cells vs. $t$ in a deme with $K=100$. Deme extinction is not observed. Transient coexistence of $W$ and $M$ is followed by the fixation of $W$ (blue traces) or $M$ (red traces). Here $s=0.01$. (c) Invasion of $W$ deme by an $M$ cell: An $M$ cell migrates to a neighbouring $W$ deme with migration rate $m$ after a mean time $\Delta t=1/(mK)$, and then either quickly fixes, producing a new $M$ deme (right), or does not fix leaving the pair of $M$ and $W$ demes unchanged (left). The same picture holds for the invasion of an $M$ deme by a $W$ cell, see text. (d) Intermediate regime (here for the clique): The dynamics is characterised by deme extinction and $W/M$ competition (see text). Deme extinction occurs after a mean time $\tau_{E}$, and empty demes are then recolonised by an invader from a neighbouring surviving deme after $\Delta t\sim 1/(mK)$. A recolonised deme is rapidly taken over (in $\Delta t\sim\mathcal{O}(1)$). (e) Coarse-grained description of the metapopulation dynamics: Each deme is always either fully $W$ (blue / dark) or $M$ (red / light) or empty (white). In this description, different scenarios arise, shown for the clique. Competition-dominated regime: all demes are occupied and there is always fixation of $W$ or $M$. Intermediate regime: eventually $W$ or $M$ take over all occupied demes, resulting in a dynamic equilibrium of empty demes and $W$ or $M$ demes, see text. Extinction-dominated regime: there are frequent deme extinctions and the metapopulation quickly goes extinct.
@@
We consider a class of spatially explicit metapopulation models of $\Omega$ demes labelled by $x\in\{1,...,\Omega\}$, each of which at time $t$ consists of a well-mixed subpopulation of $n_{W}$ cells of wild-type $W$, and $n_{M}$ mutants of strain $M$. Wild-type cells have a baseline fitness $f_{W}=1$ and mutant cells have fitness $f_{M}=1+s$. We assume $0<s\ll 1$, giving $M$ a small selective advantage over $W$.
@@
The microbial metapopulation can be envisioned as a graph whose nodes $x\in\{1,\dots,\Omega\}$ are demes (also called sites). Each deme is a well-mixed subpopulation of size $n=n_{W}+n_{M}$ located at a node of the metapopulation graph. Here, we focus on fully-connected graphs (as in the island model  [33, 34]), called cliques, and periodic one- and two-dimensional lattices called respectively cycles and grids, see Fig. 1(a). These are regular circulation graphs, generally denoted by ${\rm G}=\{{\rm clique,cycle,grid}\}$, of $\Omega$ demes connected by edges to their $q_{{\rm G}}$ nearest neighbours via cell migration at per capita rate $m$ (independently from division and death) [33, 34, 43, 41, 42, 60], see Fig. 1(a-c). We study the eco-evolutionary dynamics of the metapopulation in the biologically relevant regime of slow migration [61, 62, 41, 42, 63], see Appendix B, and consider that initially one deme is occupied entirely by mutants ($M$ deme), while the other $\Omega-1$ demes ($W$ demes) are all populated by $W$ cells, see Sec. V. All demes are assumed to have the same carrying capacity $K$ which encodes environmental variability. In Sec. III, we assume that $K$ is constant, and in Sec. IV we let the carrying capacity switch endlessly between two values representing mild and harsh conditions [16, 17, 18, 23, 20, 19, 21, 22, 24, 32], see below.
@@
In close relation to the Moran process [64, 65, 66, 67] (see Appendix B), a reference model in mathematical biology [7], the intra-deme dynamics in a deme $x$ is thus represented by a multivariate birth-death process defined by birth and death of a cell of type $\alpha\in\{W,M\}$ in that site according to the reactions [16, 17, 22, 22, 32]
@@
where $\overline{f}\equiv(n_{W}f_{W}+n_{M}f_{M})/n$ is the the average fitness in deme $x$, and $K$ here denotes the constant carrying capacity in a static environment and its time-switching version in a dynamical environment, see below. In this formulation, without loss of generality, selection operates on birth events. This can be generalised to include selection on deaths, see e.g. [10, 11].
@@
[…]
The full individual-based model is therefore a continuous-time multivariate Markov process defined by the reaction and transition rates Eqs. (1)-(4) that satisfies the master equation Eq. (27) discussed in Appendix A.1. The microscopic intra- and inter-deme dynamics encoded in the master equation (27) has been simulated using the Monte Carlo method described in Appendix I. The eco-evolutionary dynamics of a single deme is outlined in Appendix A.2. It is worth noting that $n,n_{W}$, and $n_{M}$ are all fluctuating quantities that depend on the site $x$ and time $t$, and on $\xi$ in a time-varying environment. However, for notational simplicity, we often drop the explicit dependence on some or all of the variables $x,t$, and $\xi$. Below, we combine coarse-grained analytical approximations and individual-based stochastic simulations to study how the spatial structure, migration, and demographic fluctuations influence the fixation and extinction properties of the microbial metapopulation.
@@ III Static environments
We first consider a static environment where the carrying capacity $K$ of each deme is constant. In this setting, the size of each deme rapidly reaches $K$, with $n\approx K$, see Fig. 1(b), and the expected number of migrants per unit time and deme is $mK$. The occurrence of migration events, alongside the competition between $M$ and $W$ to take over demes of the other type, increases with $K$. Cell migration and competition are however limited when $K$ is small: regardless of their type, demes of small size are prone to extinction in a mean time $\tau_{E}(K)$, see Fig. 1(b,d). For independent demes of size $K$, the deme mean extinction time $\tau_{E}(K)$ can be obtained from a logistic birth-death process, see Appendix C, yielding
@@
when $K\gg 1$, see Fig. 2(e). In our analysis, we distinguish between different dynamical scenarios through
@@
The expression of $\Omega_{\text{occ}}$ ignores spatial correlations and hence is not accurate for cycles and grids if $\psi(m,K)$ is not much larger than 1. However, $\psi(m,K)$ allows us to efficiently distinguish between the regimes dominated by $M/W$ competition, deme extinction, and the crossover intermediate regime, see Fig. 9 and Appendix D.
@@
Henceforth, we refer to “invasion” when a cell of type $M/W$ migrates to and fixates in a $W/M$ deme, and to “recolonisation” when a cell of either type migrates into an empty deme and takes it over, see Fig. 1(c,d). Accordingly, the competition-dominated and intermediate regimes are respectively characterised by invasions and recolonisations, see Fig. 1(c,d).
III.0.1 Competition-dominated dynamics
@@
Figure 2: (a-d): Competition-dominated dynamics, $\psi\gg 1$. (a) $M$ fixation probability $\phi$ vs. constant carrying capacity $K$; (b) unconditional mean fixation time $\theta$ vs. $K$; (c) $\phi$ vs. per capita migration rate $m$; (d) $\theta$ vs. $m$. Markers are simulation results and lines are predictions of Eq. (12) for $s=0.1$ (blue) and $s=0.01$ (red) on a clique (solid lines / crosses), cycle (dashed lines / circles), and grid (dotted lines / triangles). In (a,b), $m=10^{-4},\Omega=16$, and in (c,d), $K=50,\Omega=16$. In (a,c), markers for the same $s$ are almost indistinguishable indicating independence of the spatial structure. (e,f): Extinction-dominated dynamics, $\psi<1$. (e) Mean extinction time of a single deme $\tau_{E}$ vs. $K$ ($m=0$). Blue circles are simulation data, black line line shows the predictions of Eq. (6). (f) Metapopulation mean extinction time $\theta_{E}$ vs. $K$ for $\Omega=16$ and $m=10^{-2}$ (blue) and $10^{-4}$ (red). Markers are simulation results and thick lines are predictions of Eq. (39) for cliques (solid lines / crosses), cycles (dashed lines / circles), and grids (dotted lines / triangles). Thin dashed vertical lines are guides to the eye showing $\psi=1$ for $m=10^{-2}$ (blue) and $10^{-4}$ (red). Selection plays no role in this regime, so panels (e,f) have been obtained with $s=0$. In all panels, there is initially one $M$ deme and $\Omega-1$ demes occupied by $W$. Error bars are plotted in each case but are typically too small to see.
@@
When $\psi\gg 1$ and $m\ll 1$, the carrying capacity is large enough for many migrations to occur on the timescale $\tau_{E}$ of deme extinction. Since every deme expects many incoming cells in time $\tau_{E}$, deme extinction is unlikely and can be neglected. In this regime, the dynamics is dominated by local $M/W$ competition: $W$ and $M$ cells compete in each deme to fixate the local subpopulation, see Fig. 1(b).
@@
As in Refs. [41, 51, 42], we can adopt a coarse-grained description treating each deme as a single entity of type $W$ or $M$. In doing so, we assume that migration is slow so that the mean time for an $M$ or $W$ invader to fixate in a deme is negligible compared to $1/(mK)$, the expected time between migrations, see Appendix C. In this regime, each sequential migration is an invasion attempt, with a cell from an $M/W$ site trying to invade a neighbouring $W/M$ deme, see Fig. 1(c). Here, an $M/W$ invasion is the fixation of a single $M/W$ mutant in a deme consisting of $K-1$ cells of type $W/M$.
@@
In each time unit, a deme receives from and sends to its neighbours an average of $mK$ cells. Importantly, only edges connecting $M$ and $W$ demes can lead to invasions, see Fig. 1(c). These are “active edges” and their number in state $i$ on graph ${\rm G}$ is denoted by $E_{{\rm G}}(i)$, where here we consider ${\rm G}=\{{\rm clique,cycle,grid}\}$. Moreover, migration from a deme can occur to any of the $q_{\rm G}$ neighbours of the deme, where $q_{\rm clique}=\Omega-1$, $q_{\rm cycle}=2$, $q_{\rm grid}=4$, and $q_{\rm d-dim}=2d$ for a $d$-dimensional regular lattice, see Eq. (26). The number of active edges generally varies with the metapopulation state and the spatial structure, and is difficult to determine. However, a clique being the fully-connected graph, the $i$ demes of type $M$ are connected to the $\Omega-i$ demes of type $W$, yielding $E_{{\rm clique}}(i)=i(\Omega-i)$. For a cycle, if the initial state is $i=1$, the initial $M$ deme is exactly connected to two $W$ demes. This property is conserved by the coarse-grained dynamics, with an unbreakable cluster of $M$ demes always connected to a cluster of $W$ demes by two active edges until $W$ or $M$ fixes the metapopulation, yielding $E_{{\rm cycle}}(i)=2~{}\text{~{}for $i\neq 0,\Omega$}$, see Fig. 1(a,c) and below. The number of active edges in a grid is difficult to find due to complex spatial correlations, but in Appendix E, we show that its average starting from the metapopulation state $(i,\Omega-i)=(1,\Omega-1)$, can be approximated by $2\sqrt{\pi i}$. We will therefore approximate $E_{{\rm grid}}(i)\approx 2\sqrt{\pi i}$ for $i\neq 0,\Omega$. When $i=0,\Omega$, one strain fixates the entire metapopulation, where all demes are $M$ if $i=\Omega$ and all demes are $W$ when $i=0$, and hence $E_{{\rm G}}(0)=E_{{\rm G}}(\Omega)=0$.
In the coarse-grained description of the competition-dominated dynamics, starting from a single $M$ deme ($i=1$), there are $mKE_{\rm G}(i)/q_{\rm G}$ expected migration attempts to grow an unbreakable cluster of connected $M$ demes by invading neighbouring $W$ demes. Since the probability of an $M$ invasion is $\rho_{M}$, given by Eq. (9), the $M$-cluster grows at a rate $mKE_{\rm G}(i)\rho_{M}/q_{\rm G}$. Similarly a $W$ invader attempts to grow an unbreakable $W$-cluster by invading $M$ demes, hence reducing the size of the $M$ cluster, at a rate $mKE_{\rm G}(i)\rho_{W}/q_{\rm G}$. In this representation, $M$ and $W$ invasions therefore act at the interface of an unbreakable cluster of $M$ demes by increasing or reducing its size $i$ at respective rates [41, 42, 60],

### v2

@@ Abstract
Bacteria evolve in volatile environments and complex spatial structures. Migration, fluctuations and environmental variability therefore have a significant impact on the evolution of microbial populations. Here, we consider a class of spatially explicit metapopulation models arranged as regular (circulation) graphs where wild-type and mutant cells compete in a time-fluctuating environment where demes (subpopulations) are connected by slow cell migration. The carrying capacity is the same at each deme and endlessly switches between two values associated to harsh and mild environmental conditions. It is known that environmental variability can lead to population bottlenecks, following which the population is prone to fluctuation-induced extinction. Here, we analyse how slow migration, spatial structure, and fluctuations affect the phenomena of fixation and extinction on clique, cycle, and square lattice metapopulations. When the carrying capacity remains large, bottlenecks are weak and deme extinction can be ignored. The dynamics is thus captured by a coarse-grained description within which the probability and mean time of fixation are obtained analytically. This allows us to show that, in contrast to what happens in static environments, the mutant fixation probability depends on the rate of migration. We also show that the fixation probability and mean fixation time can exhibit a non-monotonic dependence on the switching rate. When the carrying capacity is small under harsh conditions, bottlenecks are strong, and the metapopulation evolution is shaped by the coupling of deme extinction and strain competition. This yields rich dynamical scenarios, among which we identify the best conditions to eradicate mutants without dooming the metapopulation to extinction. We offer an interpretation of these findings in the context of an idealised treatment strategy and discuss possible generalisations of our models.
@@ I Introduction
In the next section, we introduce the explicit spatial metapopulation model that we study and outline our methodology. In Sec. III, we present our results in the case of static environments. This paves the way to the detailed analysis in time-fluctuating environments of Sec. IV, with the weak and strong bottleneck regimes respectively studied in Secs. IV.1 and IV.2. Sec. V is dedicated to a discussion of our findings, assumptions and possible generalisations. We present our conclusions in Sec. VI. Additional technical details are given in a series of appendices.
@@ II Model & Methods
Figure 1: Metapopulation dynamics in a static environment. (a) Examples of metapopulation graphs: a clique, cycle, and grid (from left to right). Neighbouring demes are connected by migration (double arrows). Initially, there is one mutant deme (red/light) and $\Omega-1$ wild-type demes (blue / dark), and all demes have the same constant carrying capacity $K$. (b) Dynamics in a single deme. Left: Wild-type $W$ cells (blue / dark) compete with mutants of type $M$ (red / light). When $K$ is small, the deme is prone to extinction. When $K$ is large, both types coexist prior to $W$ or $M$ fixation. Top right: Realisations of the rescaled deme size $n/K$ vs. time $t$ for $K=5$ (orange/light) and $K=100$ (green/dark) illustrating how $n$ fluctuates about $K$. Bottom right: Fraction of $M$ cells vs. $t$ in a deme with $K=100$. Deme extinction is not observed. Transient coexistence of $W$ and $M$ is followed by the fixation of $W$ (blue traces) or $M$ (red traces). Here $s=0.01$. (c) Invasion of $W$ deme by an $M$ cell: Any $M$ cell of a deme migrates to a neighbouring $W$ site with migration rate $m$ after a mean time $\Delta t=1/(mK)$, and then type $M$ either quickly fixates, producing a new $M$ deme (right), or does not fix leaving the pair of $M$ and $W$ demes unchanged (left). The same picture holds for the invasion of an $M$ deme by a $W$ cell; see text. (d) Intermediate regime (here for the clique): The dynamics is characterised by deme extinction and $W/M$ competition (see text). Deme extinction occurs after a mean time $\tau_{E}$, and empty demes are then recolonised by an invader from a neighbouring surviving deme after $\Delta t\sim 1/(mK)$. A recolonised deme is rapidly taken over (in $\Delta t\sim\mathcal{O}(\ln(K))$). (e) Coarse-grained description of the metapopulation dynamics: Each deme is always either fully $W$ (blue / dark) or $M$ (red / light) or empty (white). In this description, different scenarios arise, shown for the clique. Competition-dominated regime: all demes are occupied and there is always fixation of $W$ or $M$. Intermediate regime: eventually $W$ or $M$ take over all occupied demes, resulting in a dynamic equilibrium of empty demes and $W$ or $M$ demes; see text. Extinction-dominated regime: there are frequent deme extinctions and the metapopulation quickly goes extinct.
@@
We consider a class of spatially explicit metapopulation models of $\Omega$ demes labelled by $x\in\{1,...,\Omega\}$, each of which at time $t$ consists of a well-mixed subpopulation of $n_{W}$ cells of wild-type $W$, and $n_{M}$ mutants of strain $M$. Each wild-type cell has a baseline fitness $f_{W}=1$ and all mutant cells have fitness $f_{M}=1+s$. We assume $0<s\ll 1$, giving $M$ a small selective advantage over $W$.
@@
The microbial metapopulation can be envisioned as a graph whose nodes $x\in\{1,\dots,\Omega\}$ are demes (also called sites). Each deme $x$ is a well-mixed subpopulation of size $n(x)=n_{W}(x)+n_{M}(x)$ located at a node of the metapopulation graph. Thus, $N_{W/M}\equiv\sum_{x}n_{W/M}(x)$ is the number of $W/M$ individuals across the metapopulation, and $N\equiv N_{W}+N_{M}$ is the total number of individuals in the whole graph. Here, we focus on fully-connected graphs (as in the island model  [33, 34]), called cliques, and periodic one- and two-dimensional lattices called respectively cycles and grids; see Fig. 1(a). These are regular circulation graphs, generally denoted by ${\rm G}=\{{\rm clique,cycle,grid}\}$, of $\Omega$ demes connected by edges to their $q_{{\rm G}}$ nearest neighbours via cell migration at per capita rate $m$ (independently from division and death) [33, 34, 43, 41, 42, 60]; see Fig. 1(a-c). We study the eco-evolutionary dynamics of the metapopulation in the biologically relevant regime of slow migration (whereby intra-deme dynamics occur much faster than inter-deme dynamics; see below) [61, 62, 41, 42, 63], and consider that initially one deme is occupied entirely by mutants ($M$ deme), while the other $\Omega-1$ demes ($W$ demes) are all populated by $W$ cells; see Sec. V. All demes are assumed to have the same carrying capacity $K$ which encodes environmental variability. In Sec. III, we assume that $K$ is constant, and in Sec. IV we let the carrying capacity switch endlessly between two values representing mild and harsh conditions [16, 17, 18, 23, 20, 19, 21, 22, 24, 32]; see below.
@@
In close relation to the Moran process [64, 65, 66, 67] (see Appendix B), a reference model in mathematical biology [7], the intra-deme dynamics in a deme $x$ is thus represented by a multivariate birth-death process defined by the birth and death of a cell of type $\alpha\in\{W,M\}$ in that site according to the reactions [16, 17, 22, 22, 32]
@@
where $\overline{f}\equiv(n_{W}f_{W}+n_{M}f_{M})/n$ is the average fitness in deme $x$, and $K$ here denotes the constant carrying capacity in a static environment and its time-switching version in a dynamical environment; see below. In this formulation, without loss of generality, selection operates on birth events. This can be generalised to include selection on deaths; see e.g. [10, 11].
@@
With this, the condition for a metapopulation to be a circulation is given by
$\sum_{y~\text{n.n.}~x}\sum_{\alpha}T^{m,{\rm G}}_{\alpha}(x)=\sum_{y~\text{n.n.}~x}\sum_{\alpha}T^{m,{\rm G}}_{\alpha}(y)\quad\text{for all }x,$ (5)
where $y~\text{n.n.}~x$ denotes the sum over the $q_{G}$ neighbours $y$ of the deme $x$. In our case, this simplifies considerably to give $n(x)=\sum_{y~\text{n.n.}~x}n(y)/q_{G}$ for all $x$. Intuitively, this means that for our setup, if each deme has the same population size, then the metapopulation is a circulation: there is the same incoming and outgoing migration flow at each deme. This will indeed be the case under sufficiently large carrying capacity.
@@
driven by a random telegraph process $\xi(t)\in\{-1,1\}$. The coloured dichotomous Markov noise (DMN) $\xi(t)$ switches between $\pm 1$ according to $\xi\to-\xi$ at rate $\nu$ for the symmetric DMN (see Sec. V and Appendix H for the generalisation to asymmetric switching) [68, 69, 70]. The carrying capacity, equal across demes, thus switches at a rate $\nu$ between a value $K=K_{+}$ in a mild environment (e.g. abundance of nutrients, lack of toxin) and $K=K_{-}<K_{+}$ under harsh environmental conditions (e.g. lack of nutrients, abundance of toxin) according to $K_{-}\overset{\nu}{\rightleftharpoons}K_{+}$, and thus represents random cycles of mild and harsh conditions (feast and famine cycles). The randomly time-switching $K(t)$ drives the size of each deme, and is hence responsible for the coupling of demographic fluctuations with environmental variability, an effect particularly important when there are population bottlenecks [18, 23, 19, 21, 22, 24, 32, 16, 17]; see below. Here, the DMN is at stationarity, i.e. it is initialised from its long-time distribution where instantaneous correlations become time-independent (while the autocovariance is a function of the time difference) [71, 69]. This means that the average of the DMN vanishes, $\langle\xi(t)\rangle=0$, and its autocovariance coincides with its autocorrelation reading $\langle\xi(t)\xi(t^{\prime})\rangle=e^{-2\nu|t-t^{\prime}|}$ [68, 69, 70], where $\langle\cdot\rangle$ denotes the ensemble average and $1/(2\nu)$ is the finite correlation time (with $t,t^{\prime}\to\infty$). As a consequence, the fluctuating carrying capacity is always at stationarity, with a constant average $\langle K(t)\rangle=\langle K\rangle=(K_{+}+K_{-})/2$ and an autocorrelation $\langle K(t)K(t^{\prime})\rangle=\langle K\rangle e^{-2\nu|t-t^{\prime}|}$. In our simulations with symmetric random switching, the carrying capacity is initially drawn from its stationary distribution, with $K(0)=K_{+}$ or $K(0)=K_{-}$ each with probability $1/2$; see Sec. V and Appendix H.
The full individual-based model is therefore a continuous-time multivariate Markov process defined by the reactions and transition rates Eqs. (1)-(4) that satisfies the master equation Eq. (28) discussed in Appendix A.1. The microscopic intra- and inter-deme dynamics encoded in the master equation (28) has been simulated using the Monte Carlo method described in Appendix I. The eco-evolutionary dynamics of a single deme is outlined in Appendix A.2. It is worth noting that $n,n_{W/M}$, $T_{W/M}^{\pm}$, and $T_{W/M}^{m,{\rm G}}$ are all quantities that depend on the site $x$ and time $t$, and on $\xi$ in a time-varying environment. However, for notational simplicity, we often drop the explicit dependence on some or all of the variables $x,t$, and $\xi$. Below, we combine coarse-grained analytical approximations and individual-based stochastic simulations to study how the spatial structure, migration, and demographic fluctuations influence the fixation and extinction properties of the microbial metapopulation.
@@ III Static environments
We first consider a static environment where the carrying capacity $K$ of each deme is constant. In this setting, the size $n$ of each deme rapidly reaches and fluctuates about $K$, with $n\approx K$ when $K\gg 1$; see Fig. 1(b, top right). The expected number of migrants per unit time and deme is thus $mK$. The occurrence of migration events, alongside the competition between $M$ and $W$ to take over demes of the other type, increases with $K$. Cell migration and competition are however limited when $K$ is small: regardless of their type, demes of small size are prone to extinction in a mean time $\tau_{E}(K)$; see Fig. 1(b,d). For independent demes of size $K$, the deme mean extinction time $\tau_{E}(K)$ can be obtained from a logistic birth-death process (see Appendix C) yielding
@@
when $K\gg 1$; see Fig. 2(e). We thus refer to $K$ as “small” where fixation of a deme (intra-deme dynamics) occurs on slower timescale than its extinction. Similarly, $K$ is referred to as large (i.e. $K\gg 1$) where fixation of a site occurs faster than deme extinction. In our analysis, we distinguish between different dynamical scenarios through
@@
The expression of $\Omega_{\text{occ}}$ ignores spatial correlations and hence is not accurate for cycles and grids if $\psi(m,K)$ is not much larger than 1. However, $\psi(m,K)$ allows us to efficiently distinguish between the regimes dominated by $M/W$ competition, deme extinction, and the crossover intermediate regime; see Fig. 9 and Appendix D.
@@
Henceforth, we refer to “invasion” when a cell of type $M/W$ migrates to and fixates in a $W/M$ deme, and to “recolonisation” when a cell of either type migrates into an empty deme and takes it over; see Fig. 1(c,d). Accordingly, the competition-dominated and intermediate regimes are respectively characterised by invasions and recolonisations; see Fig. 1(c,d). A summary of the timescales for the different processes and regimes considered in this work can be found in Tables 1 and 2.
III.1 Competition-dominated dynamics
@@
Figure 2: (a-d): Competition-dominated dynamics, $\psi\gg 1$. (a) $M$ fixation probability $\phi$ vs. constant carrying capacity $K$; (b) unconditional mean fixation time $\theta$ vs. $K$; (c) $\phi$ vs. per capita migration rate $m$; (d) $\theta$ vs. $m$. Markers are simulation results and lines are predictions of Eq. (14) for $s=0.1$ (blue) and $s=0.01$ (red) on a clique (solid lines / crosses), cycle (dashed lines / circles), and grid (dotted lines / triangles). In (a,b), $m=10^{-4},\Omega=16$, and in (c,d), $K=50,\Omega=16$. In (a,c), markers for the same $s$ are almost indistinguishable indicating independence of the spatial structure. (e,f): Extinction-dominated dynamics, $\psi<1$. (e) Mean extinction time of a single deme $\tau_{E}$ vs. $K$ ($m=0$). Circles are simulation data, line shows the predictions of Eq. (7). (f) Metapopulation mean extinction time $\theta_{E}$ vs. $K$ for $\Omega=16$ and $m=10^{-2}$ (blue) and $10^{-4}$ (red). Markers are simulation results and thick lines are predictions of Eq. (39) for cliques (solid lines / crosses), cycles (dashed lines / circles), and grids (dotted lines / triangles). Thin dashed vertical lines are guides to the eye showing $\psi=1$ for $m=10^{-2}$ (blue) and $10^{-4}$ (red). Selection plays no role in this regime, so panels (e,f) have been obtained with $s=0$. In all panels, there is initially one $M$ deme and $\Omega-1$ demes occupied by $W$. In panels (a,c,f), dashed lines overlap with solid lines and so are not visible. Error bars are plotted in each case but are typically too small to see.
@@
When $\psi\gg 1$ with $m\ll 1$ (slow migration; see below), the carrying capacity is large enough for many migrations to occur on the timescale $\tau_{E}$ of deme extinction, with intra-deme dynamics occurring faster than inter-deme dynamics. Since every deme expects many incoming cells in time $\tau_{E}$, deme extinction is unlikely and can be neglected. In this regime, the dynamics is dominated by local $M/W$ competition: $W$ and $M$ cells compete in each deme to fixate the local subpopulation; see Fig. 1(b).
@@
As in Refs. [41, 51, 42], the slow migration condition allows us to adopt a coarse-grained description treating each deme as a single entity of type $W$ or $M$. This is because in the regime of slow migration, the mean time for an $M$ or $W$ invader to fixate a deme is much shorter compared to $1/(mK)$, the expected time between migrations (see Appendix C, Table 1 and Eq. 12). When the dynamics is dominated by $W/M$ competition and $m\ll 1$, at coarse-grained level, each deme can thus be regarded as being either entirely occupied by $W$ or $M$ individuals; see Fig. 1(c,e). In this regime, each sequential migration is an invasion attempt, with a cell from an $M/W$ site trying to invade a neighbouring $W/M$ deme; see Fig. 1(c). Here, an $M/W$ invasion is the fixation of a single $M/W$ mutant in a deme consisting of K1K-1 cells of type $W/M$.
@@
In each time unit, a deme receives from and sends to its neighbours an average of $mK$ cells. Importantly, only edges connecting $M$ and $W$ demes can lead to invasions; see Fig. 1(c). These are “active edges” and their number in state $i$ on graph $\rm G$} is denoted by $E_{{\rm G}}(i)$, where here we consider ${\rm G}=\{{\rm clique,cycle,grid}\}$. Moreover, migration from a deme can occur to any of the $q_{\rm G}$ neighbours of the deme, where $q_{\rm clique}=\Omega-1$, $q_{\rm cycle}=2$, $q_{\rm grid}=4$, and $q_{\rm d-dim}=2d$ for a $d$-dimensional regular lattice. The number of active edges generally varies with the metapopulation state and the spatial structure, and is difficult to determine. However, a clique being the fully-connected graph, the $i$ demes of type $M$ are connected to the $\Omega-i$ demes of type $W$, yielding $E_{{\rm clique}}(i)=i(\Omega-i)$. The $M$ demes of a clique form a single unbreakable cluster since all demes are connected. For a cycle, if the initial state is $i=1$, the $M$ deme is initially connected to exactly two $W$ demes. This property is conserved by the coarse-grained dynamics on a cycle, with an unbreakable cluster of $M$ demes always connected to a cluster of $W$ demes by two active edges until $W$ or $M$ fixes the metapopulation, yielding $E_{{\rm cycle}}(i)=2~\text{~for $i\neq 0,\Omega$}$; see Fig. 1(a,c) and below. The unbreakable nature of the $M$ cluster in these two cases (and the symmetric nature of the graphs), means there is only one possible metapopulation state for a given size of $M$ cluster. This allows us to obtain the above explicit expressions for $E_{\rm clique,cycle}(i)$.
The number of active edges in a grid is difficult to find because the cluster of demes is not unbreakable on a two-dimensional lattice, and there are many possible metapopulation states for a given number of $M$ demes. However, in Appendix E, we show that the average number of active edges on a grid, starting from the metapopulation state $(i,\Omega-i)=(1,\Omega-1)$, can be approximated by $2\sqrt{\pi i}$. We will therefore approximate $E_{{\rm grid}}(i)\approx 2\sqrt{\pi i}$ for $i\neq 0,\Omega$. When $i=0,\Omega$, one strain fixates the entire metapopulation, where all demes are $M$ if $i=\Omega$ and all demes are $W$ when $i=0$, and hence $E_{{\rm G}}(0)=E_{{\rm G}}(\Omega)=0$.
In the coarse-grained description of the competition-dominated dynamics, starting from a single $M$ deme ($i=1$), there are $mKE_{\rm G}(i)/q_{\rm G}$ expected migration attempts per unit time to grow the number of $M$ demes by invading neighbouring $W$ demes. Since the probability of an $M$ invasion is $\rho_{M}$, given by Eq. (10), the number of $M$ demes grows at a rate $mKE_{\rm G}(i)\rho_{M}/q_{\rm G}$. Similarly a $W$ invader attempts to increase the number of $W$ demes by invading $M$ demes, hence reducing the number of $M$ demes, at a rate $mKE_{\rm G}(i)\rho_{W}/q_{\rm G}$. In this representation, $M$ and $W$ invasions therefore act at the interface of $M$ and $W$ demes by increasing or reducing the number $i$ of $M$ demes at respective rates [41, 42, 60],